Use an addition or subtraction formula to find the exact value of the expression.
sin(13π/12)

Answer by lwsshak3(11628)

Use an addition or subtraction formula to find the exact value of the expression.
sin(13π/12)
**
sin(13π/12)=sin(9π/12+4π/12)=sin(3π/4+π/3)
Identity: sin(s+t)=sin s cos t+cos s sin t
sin(3π/4+π/3)=sin 3π/4 cos π/3+cos 3π/4 sin π/3
=√2/2*1/2-√2/2*√3/2
=√2/4-√6/4
=(√2-√6)/4